The admission fee of a theater is $2.50 for adults and $1.25 for children. On a certain day, 700 people went to the theater for a concert and $1375 was collected. How many children and how many adults attended the concert?
Answer by Boreal(15235)

A+C=700, A=adults and C=children
2.50A+1.25C=1375
-2.5A-2.5C=-1750, multiplying top by -2.5
-1.25C=-375
C=300 children ($375)
A=400 adults ($1000)